### v1

@@ 3.1 Modelling forecast uncertainty using a scale mixture distribution
As discussed in Section 2.2, Evidential provides the ability to perform granular attribution of uncertainty to various parts of the model (e.g., Equation (5) and (6)). However, this ability comes at the cost of model complexity and the estimated epistemic uncertainty may not be reliable (as discussed in Section 2.2). We sought to propose a simpler formulation of the problem than Evidential while addressing the concerns of Bengs et al., (2023) on hierarchical models and Meinert et al., (2022) on unresolved degrees of freedom in Evidential.
@@
We propose to simplify the model by formulating the problem as a SMD [^6] Time index $t$ has been omitted for brevity and legibility. Note that variables in this section are indexed by time for each asset: $\{y_{t},\gamma_{t},\sigma_{t}^{2},\nu_{t},\alpha_{t},\beta_{t}\}$. We use the same notations in Equation (10) as Equation (3) where the symbols have the same meaning to improve comparability. (Andrews and Mallows,, 1974),
$y\sim\operatorname{N}(\gamma,\sigma^{2}\nu^{-1}),\quad\nu\sim\operatorname{Gam}(\alpha,\beta),$ (10)
where $\nu>0$ is the scaling factor, $\operatorname{Gam}$ is the Gamma distribution, and $\alpha>1$ and $\beta>0$ are the shape and scale parameters of the Gamma distribution, respectively. Our proposed formulation effectively omits the prior on $\mu$ and places a prior on $\nu$, the scaling factor of $\sigma^{2}$. We argue that uncertainty of variance can be modelled through either $\sigma^{2}$ or $\nu$. In here, $y$ is assumed to be drawn from $\operatorname{N}(\gamma,\sigma^{2}\nu^{-1})$, with mean $\gamma$ and unknown variance $\sigma^{2}\nu^{-1}$ where $\nu$ is a latent variable that introduces uncertainty into the variance of the assumed Normal distribution of $y$. This allows flexibility to inflate the variance (by minimizing $\nu$ without inflating $\sigma^{2}$) so as to capture the extremities of the distribution. Relative to Equation (7), $\sigma^{2}$ replaces $\nu$ in the parameter set when taking the SMD approach as $\sigma^{2}$ has a richer interpretation — it directly indicates the scale of the conditional data distribution. Note that in Equation (10), placing a Gamma prior on $\nu$ is equivalent to $\sigma^{-2}\sim\operatorname{Gam}(\alpha,\beta)$ as $\nu$ and $\sigma^{-2}$ are indistinguishable in $\sigma^{2}\nu^{-1}$. However, this is distinct from using a Normal-Gamma prior as there is no Normal prior on $\mu$ in Equation (10).
The marginal distribution of a Normal distribution with unknown variance (Equation (10)) is a non-standardized t-distribution (derivation is provided in Appendix A),
@@
$\mathrm{p}(y|\gamma,\sigma^{2},\alpha,\beta)$ $=\int_{\nu=0}^{\infty}\mathrm{p}_{{}_{\rm N}}(y|\gamma,\sigma^{2}\nu^{-1})\mathrm{p}_{{}_{\rm G}}(\nu|\alpha,\beta)\mathop{}\!\mathrm{d}\nu$
@@
$=\operatorname{St}\left(y;\gamma,\frac{\sigma^{2}\beta}{\alpha},2\alpha\right).$ (11)
@@
Analogous to Equation (7), the shape parameter of this marginal Student’s t-distribution is $2\alpha$. Equation (11) is similar to Equation (4) with $y$ replacing $\mu$, and can be interpreted as the Normal distribution being “stretched out” into a heavier tailed distribution due to the uncertainty in its variance. Jointly, $\bm{\zeta}=(\gamma,\sigma^{2},\alpha,\beta)$ are parameters of the SMD distribution and are outputs of the neural network. This has the effect of regularizing the mean estimate ($\gamma$) and, similar to NIG, provides the ability to handle heavy tails of the distribution that characterize asset returns.
The corresponding NLL of Equation (11) (derivation is provided in Appendix A) is,
@@
$\mathcal{L}_{\text{SMD}}(y|\bm{\zeta})$ $=\log\left[\frac{\Gamma(\alpha)}{\Gamma(\alpha+\tfrac{1}{2})}\right]+\tfrac{1}{2}\log[2\pi\sigma^{2}\beta]+(\alpha+\tfrac{1}{2})\log\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}\beta}+1\right].$ (12)
@@
Then, $\mathcal{L}_{\text{SMD}}$ is used in place of the marginal likelihood function in Equation (2), in which the neural network learns to output parameters in $\bm{\zeta}$. In Equation (10), conditional on the scaling factor $\nu$, the data is normal with variance given by the scale of the marginal t-distribution ($\tfrac{\sigma^{2}\beta}{\alpha}$). This variance gives the uncertainty of the data. Since the predictive uncertainty given by the variance of the marginal t-distribution contains both epistemic and aleatoric uncertainties, the difference between predictive and data uncertainties gives the epistemic uncertainty. This is illustrated in Equation (13) below:
@@
Predictive $\text{Predictive uncertainty}:\operatorname{Var}[y]$ $=\tfrac{\sigma^{2}\beta}{\alpha}\cdot\tfrac{2\alpha}{2\alpha-2}=\tfrac{\sigma^{2}\beta}{\alpha-1}$
@@
Recall that the result in Equation (11) can be interpreted as a Normal distribution being stretched out into a heavier tailed t-distribution when variance is unknown. Kurtosis of the t-distribution is controlled by the shape parameter ($2\alpha$). In analysing Equation (11) and (13), we argue that $\alpha$ is analogous to “virtual observations” ($\nu$) in NIG. Epistemic uncertainty $\tfrac{\sigma^{2}\beta}{\alpha(\alpha-1)}$ is smaller than aleatoric uncertainty $\tfrac{\sigma^{2}\beta}{\alpha}$ by a factor of $\tfrac{1}{\alpha-1}$, when $\alpha>2$. Thus, as $\alpha$ increases, both epistemic uncertainty and scale of the marginal t-distribution monotonically decrease. Importantly, epistemic uncertainty also drops relative to aleatoric uncertainty, as the t-distribution converges to the Normal distribution on increasing $\alpha$. This stands in contrast to the model with NIG prior (Equation (7)), where increasing evidence $\nu$ does not monotonically lead to a decrease in scale of the t-distribution.
Our proposed SMD formulation addresses some of the concerns of Meinert et al., (2022) and Bengs et al., (2023). There are essentially three free parameters in Equation (11) as $\sigma^{2}\beta$ together should be treated as one. Parameter $\beta$ is a redundant parameter as it exists as a product together with $\sigma^{2}$ in both the marginal NLL (Equation (12)) and in all three uncertainty measures (Equation (13)). Parameters $\sigma^{2}$ and $\beta$ indicate scales of the Normal and Gamma distributions, respectively. Together, they contribute to the scale of the marginal t-distribution. The number of parameters can be reduced by either reparameterizing $\sigma^{2}\beta$ as a single parameter, or by setting $\alpha=\beta$, which we consider as the more intuitive choice. SMD encapsulates several well-known distributions as special cases. According to Andrews and Mallows, (1974) and Choy and Chan, (2008), in the case of $\alpha=\beta$, then Equation (10) is a Student’s t-distribution with $2\alpha$ degrees of freedom, and is Cauchy if $\alpha=\beta=1$. If $\alpha\neq\beta$, Equation (10) gives the Pearson Type VII (PTVII) distribution which can be re-expressed as a Student’s t-distribution in Equation (11). As epistemic uncertainty is estimated by the heavy tails of the t-distribution, we can, without loss of generality, set $\alpha=\beta$ and reformulate Equation (11) as,
@@
$\mathcal{L}_{\text{PTVII}}(y|\gamma,\sigma^{2},\alpha,\alpha)$ $=\log\left[\frac{\Gamma(\alpha)}{\Gamma(\alpha+\tfrac{1}{2})}\right]+\tfrac{1}{2}\log[2\pi\sigma^{2}\alpha]+(\alpha+\tfrac{1}{2})\log\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}\alpha}+1\right].$ (15)
@@
Comparing Equation (14) to the marginal t-distribution of using a NIG prior (Equation 7), parameters of this model relate directly to parameters of the t-distribution instead of hyperparameters of the prior distribution. Hence, mitigating the concerns of Bengs et al., (2023) on hierarchical models and Meinert et al., (2022) on unresolved degrees of freedom. Thus, we argue that SMD offers an attractive trade-off between model complexity and granularity, occupying the middle ground between Ensemble (no prior) and Evidential (prior on both mean and variance).
3.2 Architecture of the neural network
@@
Figure 1: Input sequence (shaded in red) is passed into one or more LSTM layers. Output from the LSTM layers is then fed into four subnetworks of one or more fully connected layers with ReLU activation. The final layer of each subnetwork is a fully-connected layer with linear activation. Softplus is applied to $\sigma^{2},\alpha$ and $\beta$ to ensure positivity. The four outputs of the neural network are then used to compute $\hat{y}$ and $\operatorname{Var}[\hat{y}]$ (shaded in blue). Note that we can set $\alpha=\beta$, as discussed in Section 3.1. In this case, $\beta$ is dropped and the number of subnetworks reduces from four to three.
@@
For the main application of this work, UQ of financial time-series forecasts, we propose a novel architecture for the modelling of distribution parameters, as illustrated in Figure 1. To predict $\hat{y}_{t}$, time-series inputs of both returns ($r_{t-K+1},\dots,r_{t}$) and log-transformed squared returns ($\log[r_{t-K+1}^{2}],\dots,\log[r_{t}^{2}]$) are fed into one or more long short-term memory (LSTM) layers (Hochreiter and Schmidhuber,, 1997). We log-transform squared returns to reduce skewness. The LSTM layers convert each time-series into a latent representation. The latent representation is then fed into four subnetworks, where each subnetwork is comprised of one or more fully connected layers and applies non-linear transformations on the latent representation. This allows the network to model complex relationships between the parameters in $\bm{\zeta}$ and the sequence. As noted in Section 3.1, we can set $\alpha=\beta$ and reduce the number of subnetworks to three. In other words, the network architecture illustrated in Figure 1 can be modified to output three parameters: $\bm{\zeta}=(\gamma,\sigma^{2},\alpha)$. We have kept $\beta$ to be comparable to Evidential but provide empirical results in Appendix B using the UCI dataset (the same benchmark dataset used in Lakshminarayanan et al.,, 2017 and Amini et al.,, 2020, and discussed in Appendix A.1) to show that the two networks are indeed equivalent. In the following, we explore the proposed design of the architecture in detail.
Lakshminarayanan et al., (2017) and Amini et al., (2020) introduced the Gaussian and NormalInverseGamma layers as the final layer of a neural network. These final layers output parameters of the posterior distribution. Let $\bm{a}\in\operatorname{\mathbb{R}}^{H^{(I)}}$ be the input vector of the final layer with $H^{(I)}$ dimensions and $H^{(O)}$ be the dimension of the output layer. In the case of the NormalInverseGamma layer, $H^{(O)}=4$. The NormalInverseGamma layer outputs,
$\bm{\zeta}=\mathrm{O}(\bm{a};\bm{\theta})=\bm{a}^{\mkern-1.5mu\mathsf{T}}\cdot\bm{W}^{(O)}+\bm{b}^{(O)}$
[…]
Outputs of the NormalInverseGamma layer are linear transformations of a common input $\bm{a}$ (Equation (16)). We argue that this construct is too restrictive for complex applications, such as in quantifying uncertainty of financial time-series forecasts, as detailed in Section 4.1. We propose to model each of the four parameters of SMD with a fully connected subnetwork of one or more layers. This allows for a more expressive modelling of $\bm{\zeta}$, where each parameter may have complex, non-linear relationships with the input.
@@
Additionally, we enforce constraints on $\sigma^{2}>0$, $\alpha>1$ and $\beta>0$ by applying softplus transformation with a constant term, $z^{\prime}=\log(1+\exp(z))+c$, where $z\in\{\sigma^{2},\alpha,\beta\}$ and $c$ is the minimum value of the respective parameters. The transformed values constitute the final output of the network: $\bm{\zeta}^{\prime}=\{\gamma,(\sigma^{2})^{\prime},\alpha^{\prime},\beta^{\prime}\}$. In Section 4.1, we show that this modification vastly improves quantification of forecast uncertainty of financial time-series. For other network architectures, we argue that the same approach can be applied. In the case of a feedforward network, we recommend having at least one common hidden layer that reduces the input to a single latent representation. The latent representation is then passed to individual subnetworks for specialization. We argue that the common hidden layer allows information sharing across the four parameters, while having no common hidden layer (i.e., if the input is fed into the four disjoint stacks of hidden layers directly) will prevent sharing of information across the stacks.
@@
Machine learning models are typically trained using pooled dataset of historical observations. As such, they learn the average uncertainty within the historical data. However, as noted in Section 1, asset returns exhibit time-varying volatility clustering patterns. Thus, we expect predictive uncertainty to be correlated with time-varying variance of the DGP. In other words, predictive uncertainty is high when $\sigma_{t}^{2}$ of the DGP is high and the model is “surprised” by the volatility. To inform the neural network of the prevailing volatility environment, we propose to include the log of squared returns $\{\log(r_{t-K+1}^{2}),\dots,\log(r_{t}^{2})\}$ as part of the input matrix. This follows from the use of squared returns in volatility forecasting literature (Brownlees et al.,, 2011) and allows the neural network to infer the prevailing volatility environment.
@@
Model averaging, as a special case of ensembling, is a well studied statistical method for improving predictive power of estimators (Breiman,, 1996; Goodfellow et al.,, 2016), and has previously been shown to improve accuracy of financial time-series forecasting (Wong et al.,, 2021) and sequential predictions (Raftery et al.,, 2010). As accuracy of both return forecast accuracy and predictive uncertainty are important in our motivating application, we propose to incorporate model averaging to improve return forecasts at the cost of higher predictive uncertainty estimates. For an ensemble of $M$ models, we compute the ensemble forecast $\tilde{y}$ and predictive variance $\operatorname{Var}[\tilde{y}]$ as,
@@
where $\hat{y}_{i}$ and $\operatorname{Var}[\hat{y}_{i}]$ are mean and predictive variance of model $i$, respectively. In Equation (17), $\operatorname{E}[\hat{y}^{2}]>\operatorname{E}[\hat{y}]^{2}$ (by Jensen’s inequality). Thus, predictive uncertainty of the ensemble will be higher than estimated using the marginal t-distribution alone. In Section 4, we show that model averaging resulted in significant predictive performance improvement and, despite the higher uncertainty estimates, resulted in the lowest NLL.
Popular tools for modelling time-varying volatility are Autoregressive Conditional Heteroskedasticity (ARCH; Engle,, 1982) and GARCH models. GARCH, when applied to stock returns, assumes the same DGP as Equation (1). Time-varying variance $\sigma_{t}^{2}$ is modelled using an ARMA model (Box et al.,, 1994). Parameter $\mu_{t}$ can assume a fixed value (e.g., sample mean or 0) or modelled using time-series models such as ARMA (leading to the ARMA-GARCH formulation). In our proposed framework, squared returns are provided as inputs to LSTM in similar spirit to the autoregressive terms of squared returns in GARCH. However, our proposed framework also has few differences to ARMA-GARCH. A neural network offers greater flexibility in modelling and can automatically discover interaction effects between returns and volatility. For example, higher volatility is negatively correlated with future asset returns (known as the leverage effect; Cont,, 2001). By contrast, modelling of interaction effects in additive models (such as GARCH) requires explicit specification by the user. LSTM can also be interpreted as having dynamic autoregressive orders (as opposed to fixed orders in GARCH). The input and forget gates of LSTM allow the network to control the extent of long-memory depending on features of the time-series. Multi-step ahead forecasting is an iterative process for ARMA-GARCH and forecast errors may compound. LSTM is able to predict multi-step ahead directly. In Section 4.1, we apply our framework to forecast forward 1-month U.S. stock returns using daily returns. Nonetheless, we do not directly compare against ARMA-GARCH models for two reasons. First, in this work, we are focused on advancing UQ methodologies for neural networks. We argue that several of our advances can be beneficial to both time-series and non-time-series datasets (as demonstrated in Appendix A.1). Second, we lean on the plethora of literature in comparing LSTM to ARMA-variants (e.g., Siami-Namini et al.,, 2018) and ARCH-variants (e.g., Liu,, 2019).
For ease of comparison, we outline the differences of our method to Ensemble (Lakshminarayanan et al.,, 2017) and Evidential (Amini et al.,, 2020) in Table 1.
@@
Output layer Single layer $\mu,\sigma^{2}$ Single layer $\gamma,\nu,\alpha,\beta$ Multi-layer $\gamma,\sigma^{2},\alpha,\beta$

### v2

@@ 2.4.1 Scale mixture distribution
As discussed in Section 2.3.1, Evidential provides the ability to perform granular attribution of uncertainty to various parts of the model (e.g., Equation (6) and (8)). However, this ability comes at the cost of model complexity and the estimated epistemic uncertainty may not be reliable (see Section 2.3.2). We sought to propose a simpler model formulation addressing the concerns of Bengs et al., (2023) on hierarchical models with latent variables and Meinert et al., (2022) on overparametrization with unresolved degrees of freedom in the Evidential method.
@@
We propose to simplify the model by formulating the problem as a SMD [^10] Time index $t$ has been omitted for brevity and legibility. Note that variables in this section are indexed by time for each asset: $\{y_{t},\gamma_{t},\sigma_{t}^{2},\nu_{t},\alpha_{t},\beta_{t}\}$. We use the same notations in Equation (13) as Equation (4), where the symbols have the same meaning to improve comparability. (Andrews and Mallows,, 1974),
$y|\nu\sim\operatorname{N}(\gamma,\sigma^{2}\nu^{-1}),\quad\nu\sim\operatorname{Gamma}(\alpha,\beta),$ (13)
where $\nu>0$ is the scaling factor. This scaling factor affords flexibility to inflate the variance (by minimizing $\nu$ without inflating $\sigma^{2}$) so as to capture the extremities of the distribution.
We note that uncertainty of variance can be modelled through either $\sigma^{2}$ or $\nu^{-1}$ as they are indistinguishable in $\sigma^{2}\nu^{-1}$ in Equation (13). Hence, placing the IG prior on $\sigma^{2}$, that is $\sigma\sim\operatorname{IG}(\alpha,\beta)$, is equivalent to placing a Gamma prior on the scaling factor $\nu$. However, our proposed formulation is distinct from using a NIG prior as it effectively omits the prior on $\mu$ and lets $y$ to take the distribution of $\mu$. This is motivated by our asset return forecasting application, where the mean is typically close to zero (in scale of 0.01) and thus uncertainty of $\mu$ is negligible, whereas volatility is significantly larger (standard deviation in scale of 0.1). This idea is consistent with fitting a return series with models such as Generalized ARCH (GARCH; Bollerslev, 1986a, ) in which the mean process is typically assumed zero or first order autoregressive (Carroll and Kearney,, 2009).
The marginal distribution of the Normal model in Equation (13) is a non-standardized t-distribution,
@@
$\mathrm{p}(y|\gamma,\sigma^{2},\alpha,\beta)$ $=\int_{\nu=0}^{\infty}\mathrm{p}_{{}_{\operatorname{N}}}(y|\gamma,\sigma^{2}\nu^{-1})\mathrm{p}_{{}_{\operatorname{G}}}(\nu|\alpha,\beta)\mathop{}\!\mathrm{d}\nu$
@@
$=\operatorname{St}\left(y;\gamma,\frac{\sigma^{2}\beta}{\alpha},2\alpha\right),$ (14)
@@
which is similar to Equation (7) with $y$ replacing $\mu$, has the same degrees of freedom parameter and $\sigma^{2}$ replacing $\nu$ in the parameter set. This is an advantage as $\sigma^{2}$ has a richer interpretation — it directly indicates the scale of the conditional data distribution. Derivation of Equation (14) is provided in Appendix A. Conceptually, this marginal t-distribution can be interpreted as the Normal distribution being “stretched out” into a heavier tailed distribution due to the uncertainty in its variance and provides the ability to handle heavy tails of the distribution that characterize asset returns.
Jointly, $\bm{\zeta}=(\gamma,\sigma^{2},\alpha,\beta)$ are parameters of the SMD distribution and the outputs of the neural network that minimize the NLL loss function,
@@
$\mathcal{L}_{\operatorname{SMD}}(y|\bm{\zeta})$ $=\log\left[\frac{\Gamma(\alpha)}{\Gamma(\alpha+\tfrac{1}{2})}\right]+\tfrac{1}{2}\log[2\pi\sigma^{2}\beta]+(\alpha+\tfrac{1}{2})\log\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}\beta}+1\right],$ (15)
of the t-distribution in Equation (14). Derivation is provided in Appendix A. The neural network learns to output parameters in $\bm{\zeta}$ which minimize the NLL, with $\mathcal{L}_{\operatorname{SMD}}$ replacing the marginal likelihood function $\log\mathrm{p}\big{(}y_{q}|F(\bm{x}_{q};\bm{\theta}^{*})\big{)}$ in Equation (2).
@@
For the sources of uncertainty, the data $Y|\nu$ conditional on the scaling factor $\nu$ in Equation (13) is Normal with variance given by the scale ($\tfrac{\sigma^{2}\beta}{\alpha}$) of the marginal t-distribution. This variance gives the uncertainty of the data (i.e., aleatoric uncertainty). The predictive uncertainty, containing both epistemic and aleatoric uncertainties, is given by the variance of the marginal t-distribution. Hence, the difference between predictive and aleatoric uncertainties gives epistemic uncertainty. This is illustrated in Equation (17) below:
@@
Predictive $\text{Predictive uncertainty}:\operatorname{Var}[y]$ $=\tfrac{\sigma^{2}\beta}{\alpha}\cdot\tfrac{2\alpha}{2\alpha-2}=\tfrac{\sigma^{2}\beta}{\alpha-1}$ (16)
@@
We note that epistemic uncertainty $\tfrac{\sigma^{2}\beta}{\alpha(\alpha-1)}$ is smaller than aleatoric uncertainty $\tfrac{\sigma^{2}\beta}{\alpha}$ by a factor of $\tfrac{1}{\alpha-1}$, when $\alpha>2$. Thus, as $\alpha$ increases, both epistemic uncertainty and scale of the marginal t-distribution monotonically decrease. Hence, we argue that $\alpha$ is analogous to “virtual observations”, similar to $\nu$ in NIG. More importantly, epistemic uncertainty also drops relative to aleatoric uncertainty, as the t-distribution converges to the Normal distribution on increasing $\alpha$. This stands in contrast to the model with NIG prior in Equation (9), where increasing evidence $\nu$ does not monotonically lead to a decrease in scale ($\tfrac{\beta(1+\nu)}{\nu\alpha}$) of the t-distribution.
2.4.2 Addressing the shortcomings of Evidential method
Our proposed SMD formulation addresses the concerns of Bengs et al., (2023) with latent variables and Meinert et al., (2022) with overparametrization. Regarding overparametrization, SMD effectively has only three free parameters in Equation (14) as parameter $\beta$ is redundant when it appears as a product with $\sigma^{2}$ in both the marginal NLL (Equation (15)) and all three uncertainty measures (Equation (17)). Parameters $\sigma^{2}$ and $\beta$ indicate scales of the Normal and Gamma distributions, respectively. Together, they contribute to the scale of the marginal t-distribution. Hence, the number of parameters can be reduced by reparameterizing $\sigma^{2}\beta$ as a single parameter.
Moreover, one can set $\alpha=\beta$, which we consider as the more intuitive choice. SMD encapsulates several well-known distributions as special cases. In the case of $\alpha=\beta$, Equation (13) is a Student’s t-distribution with $2\alpha$ degrees of freedom, and is Cauchy if $\alpha=\beta=1$ (Andrews and Mallows,, 1974; Choy and Chan,, 2008). If $\alpha\neq\beta$, Equation (13) gives the Pearson Type VII (PTVII) distribution which can be re-expressed as a Student’s t-distribution in Equation (14). We can, without loss of generality, set $\alpha=\beta$ and reformulate Equation (14) as,
@@ 2.4.2 Addressing the shortcomings of Evidential method
$\mathcal{L}_{\text{PTVII}}(y|\gamma,\sigma^{2},\alpha)$ $=\log\left[\frac{\Gamma(\alpha)}{\Gamma(\alpha+\tfrac{1}{2})}\right]+\tfrac{1}{2}\log[2\pi\sigma^{2}\alpha]+(\alpha+\tfrac{1}{2})\log\left[\frac{(y-\gamma)^{2}}{2\sigma^{2}\alpha}+1\right].$ (19)
@@
Comparing Equation (18) to the marginal t-distribution using a NIG prior in Equation 9, parameters of SMD relate directly to parameters of the t-distribution instead of hyperparameters of the prior distribution, mitigating the concerns of Bengs et al., (2023) on hierarchical models that estimating hyperparameters $\alpha,\beta$ relies on latent variables $\mu,\sigma^{2}$. Thus, we argue that SMD offers an attractive trade-off between model complexity and granularity, occupying the middle ground between Ensemble (no prior) and Evidential (prior on both mean and variance). Lastly, despite advances in regularizing Evidential, Combined with a simplified statistical construct displays superior results in Section 3.1 without the need for regularization.
2.4.3 Architecture of the neural network
@@
Figure 2: Architecture of our proposed Combined method. Output from the LSTM layers is fed into four subnetworks of one or more blocks, each consisting of a fully connected layer and dropout. The final layer of each subnetwork is a fully-connected layer (shaded in yellow). Softplus is applied to $\sigma^{2},\alpha$ and $\beta$ to ensure positivity. The four outputs (shaded in blue) of the neural network are then used to compute $\hat{y}$ and $\operatorname{Var}[\hat{y}]$ in Equation (17). Note that we can set $\alpha=\beta$, as discussed in Section 2.4.1. In this case, $\beta$ is dropped and the number of subnetworks reduces from four to three.
The main application of this paper is the UQ of financial time-series forecasts. In addition to the statistical model structures, for UQ in financial time-series, network architectural design is also important. We propose a novel architecture with three distinct improvements. The proposed architecture is illustrated in Figure 2 and summarized below. To predict $\hat{y}_{t}$, time-series inputs of both returns ($r_{t-K+1},\dots,r_{t}$) and log-transformed squared returns ($\log[r_{t-K+1}^{2}],\dots,\log[r_{t}^{2}]$) are fed into one or more LSTM layers. We log-transform squared returns to reduce skewness. The LSTM layers convert each time-series into a latent representation. The latent representation is then fed into four independent subnetworks, where each subnetwork is comprised of one or more blocks. Each block contains a fully connected layer, ReLU and dropout. In the following, we detail our network architectural design choices.
@@
Firstly, we apply separate subnetworks to each model parameter in $\bm{\zeta}$, whereas in Evidential, the NormalInverseGamma layer derives four hyperparameters by a linear transformation of a common input $\bm{a}$ (Equation (12)). This is clearly illustrated in Figure 2 relative to Figure 1. We argue that this construct is too restrictive for complex applications, such as in quantifying uncertainty of financial time-series forecasts. We propose to model each of the four parameters of SMD with a fully connected subnetwork of one or more layers. This allows for a more expressive modelling of $\bm{\zeta}$, where each parameter may have complex, non-linear relationships with the input. As noted in Section 2.4.2, we can set $\alpha=\beta$ and reduce the number of subnetworks to three. In other words, the network architecture illustrated in Figure 2 can be modified to output three parameters: $\bm{\zeta}=(\gamma,\sigma^{2},\alpha)$. We keep $\beta$ to be comparable to Evidential method in Section 3.1 but provide empirical results of setting $\alpha=\beta$ in Appendix C using the UCI dataset (the same benchmark dataset used in Lakshminarayanan et al.,, 2017 and Amini et al.,, 2020, and discussed in Appendix B) to show that the two networks are indeed equivalent.
@@
In each subnetwork, we enforce constraints on $\sigma^{2}>0$, $\alpha>1$ and $\beta>0$ by applying softplus transformation with a constant term, $z^{\prime}=\log(1+\exp(z))+c$, where $z\in\{\sigma^{2},\alpha,\beta\}$ and $c$ is the minimum value of the respective parameters. The transformed values constitute the final output of the network: $\bm{\zeta}^{\prime}=\{\gamma,(\sigma^{2})^{\prime},\alpha^{\prime},\beta^{\prime}\}$. In Section 3.1, we show that this modification vastly improves quantification of forecast uncertainty of financial time-series. For other network architectures, we argue that the same approach can be applied. In the case of a feedforward network, we recommend having at least one common hidden layer that reduces the input to a single latent representation. The latent representation is then passed to individual subnetworks for specialization. We argue that the common hidden layer allows information sharing across the four parameters, while having no common hidden layer (i.e., if the input is fed into the four disjoint stacks of hidden layers directly) will prevent sharing of information across the stacks.
@@
Secondly, we propose to include the log of squared returns $\{\log(r_{t-K+1}^{2}),\dots,$ $\log(r_{t}^{2})\}$ as part of the input matrix. In pooled panel datasets, the model typically learns the average uncertainty within the historical data. However, as noted in Section 1.1, asset returns exhibit time-varying volatility clustering patterns in which the predictive uncertainty is expected to be correlated with time-varying variance of the DGP. Following the use of squared returns in volatility forecasting literature (Brownlees et al.,, 2011), squared returns allow the neural network to infer the prevailing volatility environment. Squared returns are fed into LSTM in similar spirit to the autoregressive terms of squared returns in Generalized Autoregressive Conditional Heteroskedasticity (GARCH; Bollerslev, 1986b, ), a popular tool for modelling time-varying volatility in statistical models. GARCH models adopt the same DGP as Equation (1) with time-varying variance $\sigma_{t}^{2}$ modelled using an ARMA model (Box et al.,, 1994) and time-varying mean $\mu_{t}$ assuming a fixed value, 0 or some time-series models such as ARMA (leading to the ARMA-GARCH formulation).
However, our proposed framework also has a few improvements to ARMA-GARCH models as neural networks offer greater flexibility in modelling and can automatically discover interaction between returns and volatility. This interaction is known as the leverage effect (Cont,, 2001), where periods of higher volatility is negatively correlated with future asset returns. By contrast, modelling of interaction effects in GARCH models requires explicit specification by the user in most packages. LSTM can also be interpreted as having dynamic autoregressive orders (as opposed to fixed orders in GARCH). The input and forget gates of LSTM allow the network to control the extent of long-memory depending on features of the time-series. Nonetheless, we do not directly compare against ARMA-GARCH models for two reasons. First, in this paper, we are focused on advancing UQ methodologies for neural networks. We argue that several of our advances can be beneficial to both time-series and non-time-series datasets (as demonstrated in Appendix B). Second, we lean on the plethora of literature in comparing LSTM to ARMA-variants (e.g., Siami-Namini et al.,, 2018) and ARCH-variants (e.g., Liu,, 2019).
@@
Lastly, we adopt model averaging, as a special case of ensembling, for improving predictive power of estimators (Breiman,, 1996; Goodfellow et al.,, 2016). It was previously shown to improve accuracy of financial time-series forecasting (Wong et al.,, 2021) and sequential predictions (Raftery et al.,, 2010). As accuracy of both return forecast and predictive uncertainty are important, we propose to incorporate model averaging to improve return forecasts at the cost of higher predictive uncertainty estimates. For an ensemble of $M$ models, we compute the ensemble forecast $\tilde{y}$ and predictive variance $\operatorname{Var}[\tilde{y}]$ as,
@@
where $\hat{y}_{i}$ and $\operatorname{Var}[\hat{y}_{i}]$ in Equation (20) are mean and predictive variance of model $i$, respectively. The higher predictive uncertainty for return forecast $\tilde{y}$ with model averaging stems from the fact that $\operatorname{E}[\hat{y}^{2}]>\operatorname{E}[\hat{y}]^{2}$ by Jensen’s inequality in Equation (20). Thus, predictive uncertainty using model averaging contains additional uncertainty than those estimated using the marginal t-distribution alone. In Section 3.2, we show that model averaging resulted in significant predictive performance improvement and, hence lowest NLL, despite the higher uncertainty estimates. Procedural-wise, this model averaging is the same as the Ensemble method (hence the name the Combined method of Ensemble and Evidential), but the aim is to improve return forecast accuracy than capturing epistemic uncertainty in Ensemble.
For ease of comparison, we outline the differences of our combined method to Ensemble (Lakshminarayanan et al.,, 2017) and Evidential (Amini et al.,, 2020) methods in Table 1.
@@
Output layer Single layer $\mu,\sigma^{2}$ Single layer $\gamma,\nu,\alpha,\beta$ Multiple subnetworks for each of $\gamma,\sigma^{2},\alpha,\beta$